OpenStudy (anonymous):

ok so =tacrtan4t- ∫4t/(1+16t2)dt i believe I need to modify this last integral further so that I my final answer has a ln component. How do I do that?

OpenStudy (mr.math):

Substitute \(1+16t^2=u\).

OpenStudy (anonymous):

ok got it. Thanks!!!

OpenStudy (mr.math):

Or rewrite the integral as, \(\frac{1}{8}\int\limits \frac{8(4t)}{1+16t^2}dt. \)